# Efficient Sampling for Pauli Measurement-Based Shadow Tomography in Direct Fidelity Estimation

### v5

@@ V.4 Fidelity with Dicke state
From Table 1, we see that the term inside the square bracket in (V.4) is zero if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}=0$ or $\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=0$, ±$\pm 1/2^{2k-2l}$ if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=1$ and |$|\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 0\pmod{2}$, and ±$\pm i/2^{2k-2l}$ if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=1$ and |$|\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 1\pmod{2}$. Meanwhile, $\text{tr}(\rho|\mathbf{j}\rangle\langle\mathbf{i}|)$ is the complex conjugate of (V.4), and the sample
$V_{c}=\left(\prod_{m\in\text{s}(\mathbf{i})-\text{s}(\mathbf{j})}\langle 0|U_{m}^{\dagger}|\hat{b}_{m}\rangle\langle\hat{b}_{m}|U_{m}|1\rangle\prod_{m\in\text{s}(\mathbf{j})-\text{s}(\mathbf{i})}\langle 1|U_{m}^{\dagger}|\hat{b}_{m}\rangle\langle\hat{b}_{m}|U_{m}|0\rangle\right.$
@@
$o_{l}=\sum_{\begin{subarray}{c}\mathbf{i}<\mathbf{j},\\ |\mathbf{i}|=|\mathbf{j}|=k,\\ |\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})|=l\end{subarray}}\sum_{\begin{subarray}{c}U_{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})}\in T\setminus\{\mathbb{I}\},\\ U_{m\notin\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})}=\mathbb{I},\\ |\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 0\pmod{2}\end{subarray}}\mathbb{E}_{\hat{b}}\left[V_{c}\cdot\binom{n}{k}^{-1}\right].$
@@
where $\hat{o}_{l}\equiv V_{c}\cdot\binom{n}{k}^{-1}\cdot c_{l}\cdot 2^{2k-2l-1}$ and ($\mathcal{F}(\{U_{m}^{\prime}\})\in T\setminus\{\mathbb{I}\}$ such that |$|\{U\in\{U_{m}^{\prime}\}\cup\mathcal{F}(\{U_{m}^{\prime}\})|U=HS^{\dagger}\}|\equiv 0\pmod{2}$. Therefore, we have $0\leq\hat{d}\leq\binom{n}{k}^{-1}$ and $-\binom{n}{k}^{-1}\cdot c_{l}\leq\hat{o}_{l}\leq\binom{n}{k}^{-1}\cdot c_{l}$.
Define
$S\equiv\frac{1}{2}+\sum_{l=\max(0,2k-n)}^{k-1}c_{l}.$
$=d+\sum_{l=\max(0,2k-n)}^{k-1}o_{l}$
$=d+o$
$=F(\rho,\sigma).$
We generate
samples of $\hat{F}$. From Hoeffding’s inequality,
The total number of measurements is $O(n^{2k})$, where the asymptotics are with respect to $n$, while $k$, $\epsilon$, and $\delta$ remain constant. Alg. 3 provides a complete pseudocode for the Dicke state with $k$ excitations.
Table 4: ALG. 3. DFE for $n$-qubit Dicke state with $k$ excitations.
Input: Unknown state $\sigma$, error parameters $(\epsilon,\delta)$
Output: Estimate of the fidelity $F(\rho,\sigma)$, where $\rho$ is the $n$-qubit Dicke state with $k$ excitations
Define $c_{l}$ as in (16).
$\texttt{sum}\leftarrow 0$
for    in range($N$):
   if $l=k$:
    $\hat{b}=\texttt{measureComputationalBasis}(\sigma)$
[…]
   else
    $U\leftarrow\bigotimes_{m=1}^{n}U_{m}$
    $\sigma^{\prime}\leftarrow U\sigma U^{\dagger}$
    $\hat{b}=\texttt{measureComputationalBasis}(\sigma^{\prime})$
            + |$|\{m\in\text{s}(\mathbf{j})-\text{s}(\mathbf{i})\mid\hat{b}_{m}=1\}|$
[…]
   end if
end for
return $\texttt{sum}/N+1/2\binom{n}{k}$
VI Numerical simulations
Using the previous importance sampling technique, the average number of samplesand falls strictly below the bound (2) for many configurations. Furthermore, the bound (2) itself may be quite rough, as it involves two layers of estimation: one for importance sampling and another for characteristic function estimation. Thus, it is essential to experimentally demonstrate the improvement of our method over the importance sampling approach.
In practice, the precision of estimates in Algs 1 and 2 is not directly controllable, but the number of samples can be specified. Therefore, we compare Ref. [14] with Algs 1 and 2 as follows: First, Ref. [14] with parameters $(\epsilon,\delta)$ is used to estimate the fidelity. Next, Algs 1 and 2 are executed using the same number of samples as Ref. [14]. Finally, the mean absolute errors of the estimates are compared.
Input: Target state $\rho$, fidelity $f$
Output: Random state $\sigma$ such that $F(\rho,\sigma)=f$
$\sigma$ = randomDensityMatrix()
$P\leftarrow I-\rho$
$\sigma\leftarrow P\sigma P$
$\sigma\,\texttt{*=}\,(1-f)/\text{tr}(\sigma)$
return $f\rho+\sigma$
Figure 2: FIG. 2. MSE comparison for (a) GHZ state with $(\epsilon,\delta)=(0.1,0.1)$, (b) GHZ state with $(\epsilon,\delta)=(0.05,0.05)$, (c) W state with $(\epsilon,\delta)=(0.1,0.1)$, and (d) W state with $(\epsilon,\delta)=(0.05,0.05)$. The label ‘PL’ refers to the algorithm in Ref. [14], which is based on Pauli-Liouville coefficients [35]. Error bars represent 95% confidence intervals.
The states were randomly generated with fidelities ranging from 0 to 1 in increments of 0.01 (i.e., $0,0.01,0.02,\ldots,1$), as described in Alg. 4. Each state is generated by first creating a random mixed state and projecting it onto the subspace orthogonal to the target state (GHZ or W). The output state with the desired fidelity is then obtained as a linear combination of the target state and the projected state. FIG. 2 shows the comparison of mean squared error (MSE) for up to 7 qubits. Algorithms 1 and 2 significantly improve estimation accuracy over Ref. [14]. The simulations were performed using PennyLane’s classical simulation framework [36].
VII Discussion
Acknowledgements.
Author Contributions

### v6

@@ Appendix B Off-diagonal terms for Dicke state
From Table 1, we see that the term inside the square bracket in (B) is zero if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}=0$ or $\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=0$, ±$\pm 1/2^{2k-2l}$ if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=1$ and |$|\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 0\pmod{2}$, and ±$\pm i/2^{2k-2l}$ if $\delta_{\hat{b}_{m\in\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})},\mathbf{1}}\delta_{\hat{b}_{m\notin\text{s}(\mathbf{i})\cup\text{s}(\mathbf{j})},\mathbf{0}}=1$ and |$|\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 1\pmod{2}$. Meanwhile, $\text{tr}(\rho|\mathbf{j}\rangle\langle\mathbf{i}|)$ is the complex conjugate of (B), and the sample
$V_{\text{D}}=\left(\prod_{m\in\text{s}(\mathbf{i})-\text{s}(\mathbf{j})}\langle 0|U_{m}^{\dagger}|\hat{b}_{m}\rangle\langle\hat{b}_{m}|U_{m}|1\rangle\prod_{m\in\text{s}(\mathbf{j})-\text{s}(\mathbf{i})}\langle 1|U_{m}^{\dagger}|\hat{b}_{m}\rangle\langle\hat{b}_{m}|U_{m}|0\rangle\right.$
@@
$o_{l}=\sum_{\begin{subarray}{c}\mathbf{i}<\mathbf{j},\\ |\mathbf{i}|=|\mathbf{j}|=k,\\ |\text{s}(\mathbf{i})\cap\text{s}(\mathbf{j})|=l\end{subarray}}\sum_{\begin{subarray}{c}U_{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})}\in T\setminus\{\mathbb{I}\},\\ U_{m\notin\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})}=\mathbb{I},\\ |\{m\in\text{s}(\mathbf{i})\Delta\text{s}(\mathbf{j})|U_{m}=HS^{\dagger}\}|\equiv 0\pmod{2}\end{subarray}}\mathbb{E}_{\hat{b}}\left[V_{\text{D}}\cdot\binom{n}{k}^{-1}\right].$
@@
where $\hat{o}_{l}\equiv V_{\text{D}}\cdot\binom{n}{k}^{-1}\cdot c_{l}\cdot 2^{2k-2l-1}$ and ($\mathcal{F}(\{U_{m}^{\prime}\})\in T\setminus\{\mathbb{I}\}$ such that |$|\{U\in\{U_{m}^{\prime}\}\cup\mathcal{F}(\{U_{m}^{\prime}\})|U=HS^{\dagger}\}|\equiv 0\pmod{2}$. Therefore, we have $-\binom{n}{k}^{-1}\cdot c_{l}\leq\hat{o}_{l}\leq\binom{n}{k}^{-1}\cdot c_{l}$.